OpenStudy (anonymous):

2 8x10-sized portraits, 2 5x7-sized portraits, and 8 wallet-sizes portraits, what fraction of the portraits are 5x7-sized?

OpenStudy (anonymous):

1/6 is the answer because 2+2+8=12 total pictures so the 2 5X7 size are 2 out of 12 total pictures or 2/12=1/6 when reduced

OpenStudy (anonymous):

thanks